how do you solve the problem:
f(x) =1 - 1 divided by (x+2)squared and the value of x is 0

Answer by edjones(7311)

f(x)=1-(1/(x+2)^2)
f(0)=1-(1/(0+2)^2
=1-(1/4)
=3/4
.
Ed